How do you use derivatives and linear approximation (the formula for linear approximation being
L(x)=D(1-(x/60)) at the point x=0) to approximate the amount of time it takes to drive 45 miles at 62mph?

Answer by Alan3354(69443)

How do you use derivatives and linear approximation (the formula for linear approximation being
L(x)=D(1-(x/60)) at the point x=0) to approximate the amount of time it takes to drive 45 miles at 62mph?
====================
I don't.
Just use d= r*t
t = d/r